OpenStudy (anonymous):

Solution 3n / 12 +8 =13

OpenStudy (anonymous):

Rearrange and solve for n.

OpenStudy (anonymous):

3n / 12 +8 =13 3n/12=5 3n=5x12 n=60/3 n=20